i. find limπ‘₯β†’2 π‘₯^ 2βˆ’3π‘₯ +2 / π‘₯ ^2βˆ’4


ii. Where are is the function 𝑓(π‘₯) = 1 2π‘₯ ^2βˆ’6π‘₯+4 continuous?


iii. Given that given π‘Œ = tanβˆ’1 π‘₯ show that 𝑑𝑦 /𝑑π‘₯ = 1 1+π‘₯ ^2Β 



iv. Prove using the first principle that the derivative of sin π‘₯ is cos π‘₯ and that the derivative of π‘π‘œπ‘ π‘₯ is – 𝑠𝑖𝑛x


v. Differentiate sin 𝑦 βˆ’ π‘₯ 2𝑦 3 βˆ’ π‘π‘œπ‘ π‘₯ = 3y


vi. limπ‘₯β†’1 π‘₯^2+π‘₯ 3βˆ’5π‘₯+3/ π‘₯^3+2π‘₯^2+7π‘₯+4



vii. limπ‘₯β†’βˆž 2π‘₯/ 3π‘₯^6+π‘₯+4


1
Expert's answer
2021-04-15T06:31:35-0400

i.

"\lim\limits_{x\to 2}\dfrac{x^2-3x+2}{x^2-4}=\lim\limits_{x\to 2}\dfrac{(x-1)(x-2)}{(x-2)(x+2)}"

"=\lim\limits_{x\to 2}\dfrac{x-1}{x+2}=\dfrac{2-1}{2+2}=\dfrac{1}{4}"

ii. The function "f(x)=12x^2-6x+4" is continuous for all "x\in \R" as polynomial.


iii. Let "y=\tan^{-1}(x)." Then "\tan y=x, -\dfrac{\pi}{2}<y<\dfrac{\pi}{2}"



"(\tan^{-1}(x))'=\dfrac{1}{(\tan y)'}=\dfrac{1}{\dfrac{1}{\cos^2y}}"

"=\dfrac{1}{1+\tan^2y}=\dfrac{1}{1+x^2}"

iv.


"(\sin x)'=\lim\limits_{h\to 0}\dfrac{\sin(x+h)-\sin x}{h}"

"=\lim\limits_{h\to 0}\dfrac{\sin x\cos h+\cos x \sin h-\sin x}{h}"

"=\lim\limits_{h\to 0}\dfrac{\sin x(\cos h-1)+\cos x \sin h}{h}"

"=\sin x\lim\limits_{h\to 0}\dfrac{\cos h-1}{h}+\cos x\lim\limits_{h\to 0}\dfrac{\sin h}{h}"

"=\sin x(0)+\cos x(1)=\cos x"



"(\cos x)'=\lim\limits_{h\to 0}\dfrac{\cos(x+h)-\cos x}{h}"

"=\lim\limits_{h\to 0}\dfrac{\cos x\cos h-\sin x \sin h-\cos x}{h}"

"=\lim\limits_{h\to 0}\dfrac{\cos x(\cos h-1)-\sin x \sin h}{h}"

"=\cos x\lim\limits_{h\to 0}\dfrac{\cos h-1}{h}-\sin x\lim\limits_{h\to 0}\dfrac{\sin h}{h}"

"=\cos x(0)-\sin x(1)=-\sin x"

v.


"\dfrac{d}{dx}(\sin y-x^2y^3-\cos x)=\dfrac{d}{dx}(3y)"

Use the Chain Rule


"\cos y \dfrac{dy}{dx}-2xy^3-3x^2 y^2 \dfrac{dy}{dx}+\sin x=3\dfrac{dy}{dx}"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=\dfrac{-2xy^3+\sin x}{3+3x^2y^2-\cos y}"


vi.

"\lim\limits_{x\to 1}\dfrac{x^2+x^3-5x+3}{x^3+2x^2+7x+4}=\dfrac{1^2+1^3-5(1)+3}{1^3+2(1)^2+7(1)+4}"

"=0"

vii.


"\lim\limits_{x\to \infin}\dfrac{2x}{3x^6+x+4}=\lim\limits_{x\to \infin}\dfrac{\dfrac{2x}{x^6}}{\dfrac{3x^6}{x^6}+\dfrac{x}{x^6}+\dfrac{4}{x^6}}"

"=\lim\limits_{x\to \infin}\dfrac{\dfrac{2}{x^5}}{3+\dfrac{1}{x^5}+\dfrac{4}{x^6}}=\dfrac{0}{3+0+0}=0"